Graph these equations
4x+5y=-2
4y-x=11

Answer by stanbon(75887)

Graph these equations
4x+5y=-2
4y-x=11
----------------
Put both in slope-intercept form:
y = (-4/5)x - (2/5)
y = (1/4)x + (11/4)
------------------------

============================
Cheers,
Stan H.